according to a may 2013 gallup poll, 53% of americans support the legalization of same sex marriage.
b. if a random sample of 10 americans is found, how many people would we expect to support the legalization of same sex marriage
c. what would the standard deviation of this 10 person distribution be
d. if we would randomly sample 10 people, what is the probability that exactly 6 would support the legalization of same sex marriage.
e if we would randomly sample 10 people, what is the probability less than 6 would support the legalization of same sex marriage.
f. if we secure a simple randome sample of 250 americans, what is the probability that less than 120 would support the legalization of same sex marriage

Found 2 solutions by ewatrrr, stanbon:
Answer by ewatrrr(24785)

 
Hi,
b. 10*.53 = 5.3 Expect 5
c. sqrt(.53*.47/10)
d. p = .53 , n = 10 P(x=6)= binompdf(10, .53, 6)= .2271
e. P( x < 6) = binomcdf(10, .53, 5) = .5474
f. P( x < 120) = binomcdf(250, .53, 119) = .4987
recommend stattrek.com as a reference

Answer by stanbon(75887)

according to a may 2013 gallup poll, 53% of americans support the legalization of same sex marriage.
b. if a random sample of 10 americans is found, how many people would we expect to support the legalization of same sex marriage
E(x) = np = 10*0.53 = 5.3
--------------------------------------------
c. what would the standard deviation of this 10 person distribution be
std = sqrt(npq) = sqrt(5.3*0.47) = 1.58
----------------------------------------------
d. if we would randomly sample 10 people, what is the probability that exactly 6 would support the legalization of same sex marriage.
P(x = 6) = 10C6(0.53)^6(0.47)^4 = binompdf(10,0.53,6) = 0.2271
---------------------------
e if we would randomly sample 10 people, what is the probability less than 6 would support the legalization of same sex marriage.
---
P(0<= x <=5) = binomcdf(10,0.53,5) = 0.5474
--------
f. if we secure a simple randome sample of 250 americans, what is the probability that less than 120 would support the legalization of same sex marriage
P(x < 120) = binomcdf(250,0.53,119) = 0.0499
========================
Cheers,
Stan H.
========================